Galaxy Inc. has issued and outstanding a total of 30,000 shares of $3 preferred shares and 185,000 common shares. The company began operations and issued both classes of shares on January 1, 2012. It has declared and paid dividends each year as shown below.
Required:Using the table below, calculate the total dividends distributed to each class of shareholder, assuming:
a. The preferred shares are cumulative.
b. The preferred shares are non-cumulative
1
Expert's answer
2020-02-25T10:13:09-0500


• Class A with a nominal value of $ 3. and cumulative dividends at a rate of 8% per annum;

• class B with a nominal value of $ 3 and non-cumulative dividends at a rate of 12% per annum.

The corporation’s dividend policy establishes a quarterly dividend payment.

According to the results of the first, second and third quarters, the board of directors decided not to pay dividends, but in the fourth quarter it allocated $ 10,000 for dividends. At the same time, the number of shares outstanding at the end of the fourth quarter was:

Cumulative privileged: 15.000

Non-cumulative privileged: 15.000

Simple: 185.000

Assume that at the beginning of the first quarter there was no debt to pay cumulative dividends on preferred shares of class A. Although no dividends were paid in the first quarter, they were accrued on class preferred shares based on a quarterly rate of 2% (8% / 4) in size.

DI = 15.000 * 3 * 0.02 = $ 900

Similarly, dividends will be accrued in the second and third quarters.

DII = 15.000 * 3 * 0.02 = $ 900

DIII = 15.000 * 3 * 0.02 = $ 900

Thus, at the end of the third quarter, the corporation will accumulate a debt of cumulative dividends to holders of class A preferred shares in the amount of $ 2700

Since the board of directors decided to pay dividends in the fourth quarter, the corporation, first of all, is obliged to pay back the accumulated debt to the owners of class A preferred shares in the amount of $ 2700 e. In addition, their owners will also be paid dividends for the 4th quarter in the amount of $ 900

DIV = 15.000 * 3 * 0.02 = $ 900

Thus, $ 3.600 or $ 0.24 per share ($ 3.600/15.000) will be allocated for the payment of dividends on class A preferred shares.

For class B preferred shares and ordinary shares, dividends are non-cumulative in nature, therefore, no accumulation was made in the first 3 quarters. However, holders of Class B preferred shares have a preferential right to receive dividends over holders of ordinary shares. The dividend per class B share, based on a quarterly rate of 3% (12%/4), will be $ 0.09, and $ 1350 will be paid on all shares. (0.09 * 15000).

The remaining amount of $ 5050. (10.000-3600-1350) will be distributed among the holders of ordinary shares, which is $ 0.027 per share. (5050/185000).

Thus, holders of class A preferred shares will receive cumulative dividends of $ 0.24. per share, for class B their size will be $ 0.09, and $ 0.027 per ordinary share.